Jimmyjohnjoe:

Find the solution for the given equation and check the answer 4x – 7 (2-x) = 3x + 2

Jimmyjohnjoe:

@thekayla

Jimmyjohnjoe:

@saralance

Jimmyjohnjoe:

@thebatman

Jimmyjohnjoe:

@thewraith2021

Jimmyjohnjoe:

help plz

TheKayla:

@jimmyjohnjoe wrote:
help plz
The given is as follows: 4x – 7 (2-x) = 3x + 2 Step two you wanna simplify the left side of equation within the bracket like so: 4x – 14 + 7x = 3x + 2 11x – 14 = 3x + 2 After that now you want to bring all the ‘x’ terms on one side and constant the term on other side of your equation like so: 11x-3x = 14 + 2 8x = 16 X = 16/8 X = 2 After you complete those steps you should come up with the solution for the given equation 4x – 7 (2-x) = 3x + 2 which if you did it correctly should be: 2

Jimmyjohnjoe:

thx baby

Jimmyjohnjoe:

how about for this one? 4x=72

Jimmyjohnjoe:

Sandwich u didnt show ur work

Jimmyjohnjoe:

Saucein Sauce screw u Sauce

TheWraith2021:

That doesn't make any kind of sense, whatsoever.

Jimmyjohnjoe:

wsp wraith?

TheWraith2021:

Uh, hey.

Jimmyjohnjoe:

ay dis my broski wraith everybody leave him df alone

Jimmyjohnjoe:

@baddestgirl hey sexy

TheKayla:

@jimmyjohnjoe wrote:
how about for this one? 4x=72
Just divide both sides of the equation with 4: \displaystyle \frac{4x}{4}=\frac{72}{4} 4 4x ​ = 4 72 ​ \displaystyle x=18x=18

TheKayla:

It's that simple smh-

TheKayla:

1 attachment
TheKayla:

There's a calculator right under your post 🖖🙄

TheKayla:

May I get my medal now?

TheWraith2021:

He's offline.

TheKayla:

😩